The accompanying table represents the quantity produced, the total revenue, and the total cost of a firm operating in a perfectly competitive market. Refer to this table to answer the following questions.

Quantity Total Revenue Total Cost
0 $0 $3
1 $5 $5
2 $10 $9
3 $15 $13
4 $20 $19



Assuming that all firms have the same cost structure, the price is

options:

$5.

$3.

$10.

$2.

$9.
Answer #1

$5.

Reason-
Price is determined at the point at which Marginal revenue (MR) = Marginal cost (MC)
MR = Change in total revenue/Change in quantity and MC = Change in total cost/Change in quantity

When Q = 1; MR = (5-0)/(1-0) = 5 and MC = (5-3)/(1-0) = 2
When Q = 2; MR = (10-5)/(2-1) = 5 and MC = (9-5)/(2-1) = 4
When Q = 3; MR = (15-10)/(3-2) = 5 and MC = (13-9)/(3-2) = 4
When Q = 4; MR = (20-15)/(4-3) = 5 and MC = (19-13)/(4-3) = 6

Output produced is that when MR is greater than or equal to MC. We can see that MR exceeds MC upto Q = 3. So, Q = 3 is produced. Total revenue = price*quantity = 15
So, Price = 15/quantity = 15/3 = 5